The perimeter of a rectangle is 46 yd , and the area of the rectangle is 90 yd squared 2 . Find the dimensions of the rectangle.






Answer by tquon(6)

the perimeter equation would be 2w+2h=p, with w=width and h=hieght. the area equation would be w*h=a. so if 2w+2h=46 and w*h=90, you can simplify the perimeter equation to be 2w=46-2h and again to w=23-h. Then you can plug in that w for the w in the area equation. So (23-h)*h=90. Then you multiply it out to get 23h-h^2=90. subtract 90 from each side to get -h^2+23h-90=0. you can then divide by -1 on each side to make it easier, getting h^2-23h+90=0. this can be factored out to (h-18)(h-5)=0 so your possible answers for the height are 18 and 5. putting these back into the original area equation show that the dimensions are 5x18. So your answer is 5x18.